Prove that 3[sin^4(3π/2-alpha)+sin^4(3π+alpha)]-2[sin^6(π/2+alpha)+sin^6(5π-alpha)] =1. Pl send fast

Grade:10

1 Answers

Jayanth Kasaraneni
7 years ago
=3[sin^4(3π/2-alpha)+sin^4(3π+alpha)]-2[sin^6(π/2+alpha)+sin^6(5π-alpha)]=3(cos^{4}(alpha)+sin^{4}(alpha))-2(cos^{6}(alpha)+sin^{6}(alpha))
=3(1-2sin^{2}(alpha)cos^{2}(alpha))-2(1-3sin^{2}(alpha)cos^{2}(alpha))
=3-2=1